OpenStudy (anonymous):

Liner equation with 2 variable question Is this ordered pair a solution of this given equation? (8,7), y=-1/4x +5

OpenStudy (kewlgeek555):

Okay. What you have to do is plug in the numbers. Remember that (x, y), so x = 8 and y = 7.

OpenStudy (kewlgeek555):

Your equation would be: \[7=1\div4*8+5\]

OpenStudy (kewlgeek555):

Do you need help solving this equation?

OpenStudy (anonymous):

Yes thanks

OpenStudy (kewlgeek555):

Okay. Forget about the "=7" for now. Remember PEMDAS. Parenthesis, then Exponents, then Multiplication OR Division, then Addition OR Subtraction. So we have no parenthesis or exponents. We do have multiplication and division, though. Since we work on equations left to right and it is division OR multiplication, lets do division first: What is:\[1\div4\]?

OpenStudy (kewlgeek555):

Think of one whole pizza divided into four. Think about quarters.

OpenStudy (kewlgeek555):

1 divided into 4 is 0.25

OpenStudy (kewlgeek555):

Okay, so now we have a simpler equation:\[7=0.25\times8+5\]

OpenStudy (kewlgeek555):

Okay, based on PEMDAS, we multiply next. 0.25 * 8 = ?

OpenStudy (kewlgeek555):

If you don't have a calculator with you. Think about how many quarters make 1 whole.

OpenStudy (anonymous):

0.25*8=2

OpenStudy (kewlgeek555):

Great job! Now you have a simpler equation:\[7 = 2+5\] Now, I am sure you know what 2 + 5 is. It's 7. Does 7 = 7? Yes it does! Is this order pair a solution to this equation? Yes!

OpenStudy (anonymous):

Yes its a solution

OpenStudy (kewlgeek555):

So, your final answer is: "Yes, this ordered pair, (8, 7), is a solution this equation because 7 = 7."

OpenStudy (kewlgeek555):

It was a pleasure working with you! :-) Thanks for the medal.

OpenStudy (anonymous):

@kewlgeek555 I don't think that (8, 7) is a solution. You forgot to consider the negative sign (-1/4x)

OpenStudy (kewlgeek555):

:-O @theANNEbeatable You are right! :-OOO T_T

OpenStudy (kewlgeek555):

\[7=-1\div4 \times8+5\]\[7=-0.25\times8+5\]\[7=-2+5\]\[7=3\]Okay, so ... No, the ordered pair does not give a solution to this equation. Sorry @shakur

OpenStudy (kewlgeek555):

Thanks @theANNEbeatable